OpenStudy (anonymous):

How to convert (365) base 8 to the base 10.

ganeshie8 (ganeshie8):

\((365)_8\) is a short hand notation for saying you have 5 of 8^0, 6 of 8^1 3 of 8^2

ganeshie8 (ganeshie8):

\[\large (\color{red}{365})_8 = \color{Red}{5}\times 8^0 + \color{red}{6}\times 8^1 + \color{red}{3}\times 8^2\] simplify

OpenStudy (anonymous):

Okay, I got it now. Thanks ganeshie...

ganeshie8 (ganeshie8):

yw!